OpenStudy (anonymous):

hight of a window is 5√2 in. +5√2 in or 10√2 in. the length is 4(5√2) in. or 20√2 in ....what is the perimeter in inches?????

OpenStudy (anonymous):

Perimeter = 2*(height + length) = \[2*(10\sqrt{2}+20\sqrt{2}) = 60\sqrt{2}\]